OpenStudy (makayla52):

the school is selling adult and children^s tickets for play. each child ticket is $3.50 and each adult is $8.75. if 1950 total tickets were solid and the school and the school collected $13.387.50, how many of each tickets did the school sel

OpenStudy (makayla52):

I have my variables already

OpenStudy (vuriffy):

3.50x + 8.75y = 13,387.50 x + y = 1950

OpenStudy (vuriffy):

700*3.5 = 2450 1250 * 8.75 = 10,937.5 10,937.5 + 2450 = ?

OpenStudy (vuriffy):

700 + 1250 = 1950. 3.50(700) + 8.75(1250) = 13,387.50